You had $50,000 to invest. the two funds you decided to invest in are the You-Risk-It Fund that yields 14% interest and the extra-dull fund that yields 6% interest. you earned $4,500 in interest income this year. How much did you invest in each fund?
Answer by jorel1380(3719)

r+s=50000
.14r+.06s=4500
14r+6s=450000
6r+6s=300000
8r=150000
r=18750
s=31250
You invested $18750 in the You-Risk-It fund, and $31250 in the extra-dull fund..